OpenStudy (anonymous):

cot (theta) = -1/10. Where theta lies in Quadrant 2. What is csc?

OpenStudy (anonymous):

is that cosec?

OpenStudy (anonymous):

use cos2 + sin2 = 1 find sin invert it u get cosec

OpenStudy (anonymous):

cosec? you mean cos?

OpenStudy (anonymous):

wat's csc?

OpenStudy (anonymous):

What do you mean by csc? (1/sin?)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Since: \[1 + \cot ^{2} x = cosec^{2} x\] \[cosec ^{2} x = 1 - (1/10) = 9/10\] Then, I'm not exactly sure what to do, but my guess is that your root the answer.